OpenStudy (kathydimieri12):

Jake has a bag of 50 beads, some of which are blue and the remaining are green. Jake randomly pulls out a bead from the bag, records the color, and replaces it in the bag. Jake has already recorded 6 blue and 14 green beads. Based on these results, what is most likely the number of blue beads in the bag? 15 20 30 35

OpenStudy (sshayer):

50-14=?

OpenStudy (kathydimieri12):

36 @sshayer

OpenStudy (kathydimieri12):

so the answer would be D.35

OpenStudy (rayep12):

no its not

OpenStudy (rayep12):

The total population is 50 beads in the bag. The estimated percentage of blue beads is 6/20 since Jake pulled 6 blue beads out of 20 (14 + 6) attempts in his sample. If B = event of picking a blue bead with replacement, the estimated probability of picking a blue bead is: P(B) = 6/20 = 0.3. That probability is the same for each bead Jake pulls out because he replaces it each time. The best estimate of the number of blue beads in the bag is 50 x 0.3 = 15 blue beads